# Video-Language Critic: Transferable Reward Functions for Language-Conditioned Robotics

### v2

@@ B.1 Reward training
B.2 Policy training
For RL training experiments, we adapt the SAC implementation of CleanRL [44]. Policy evaluation is done every 20,000 timesteps for 50 episodes. Both the actor and critic networks contain three hidden layers of size 400, and optimization is done using Adam [45]. Other algorithm hyperparameters were kept at the implementation’s default values.
@@ B.2 Policy training
As reward model predictions for the starting state can vary across initial states, even for the same task, we shift the rewards of the episode such that $r_{1}=0$, as the subsequent behavior should be scored relative to this state. In addition to this, we apply the standard normalization logic from Gymnasium [46], which scales reward values such that their exponential moving average has fixed variance $(1-\gamma)^{2}$.
@@
Training length and hence computational cost varies considerably across tasks. We terminate training after convergence to $\geq$98success (averaged over the 10 most recent evaluations), or after a maximal training length set per task. The resulting average training length across all 25 tasks was 600k environment steps for VLC MW50 and 750k for the Open X trained model. The corresponding GPU hours vary slightly based on exact architecture used, but we obtained approximately 750k steps in 24 hours on a single NVIDIA V100 or P100 GPU. Our total computational budget was therefore 19.2 GPU hours $\times$ 5 random seeds $\times$ 25 tasks = 2,400 GPU hours for the RL training experiments with VLC MW50 and 24 $\times$ 5 $\times$ 25 = 3,000 hours with VLC Open X. For the MW40 task generalization experiments, the average training length was again 600k for VLC, but 10 random seeds were used, so the corresponding cost for training on 6 tasks was approximately 1150 GPU hours for our method, and slightly more for baselines that took longer to train.
Appendix C Meta-World environment definitions
@@
The Meta-World environments are a set of continuous control robotic manipulation tasks. The state space of each environment is $\mathbb{R}^{39}$, containing the positions and orientations of relevant objects as well as the xyz-position of the robot’s end-effector. An action $a$ in $[-1,1]^{4}$ consists of the desired end-effector xyz-translation, and one dimension for controlling the openness of the gripper. We refer to Yu et al. [11] for more details on the state information.
Appendix D Baseline implementations
@@
For the LOReL baseline experiment, we used the proposed binary classification objective, reversed negatives and 2-frame conditioning (first and last), while keeping all of our other training details and data augmentations identical to our method. This is to ensure LOReL’s smaller and older original architecture did not account for any difference in performance. The original implementations were used for all other methods, and pretrained checkpoints (e.g., HowTo100M [47] pretraining for RoboCLIP) were reused when applicable.
@@
Although not considered as a source of data in the original works, we also use our failure videos as additional negatives in the contrastive objective of LIV, RoboCLIP and R3M. Model selection was performed based on validation loss on held-out trajectories. Otherwise, the training procedure and hyperparameter settings were kept unchanged from the original work. In policy training, we apply identical reward normalization (offset and scale, as described in Section 4) for all methods, and reuse the multiplier 50 for the sparse reward.
Appendix E Robustness to visual variations
@@
Appendix F Qualitative evaluation
@@
Appendix G Model-based evaluation results per task
VLC 100 20 96 100 40 94 54 82 100 100 97 100 80
RoboCLIP 88 60 58 4 86 32 4 14 26 82 44 50 46
LOReL 26 14 44 14 4 22 22 28 72 54 100 30 36

### v3

@@ C.1 Reward training
C.2 Policy training
For RL training experiments, we adapt the SAC implementation of CleanRL (Huang et al., 2022). Policy evaluation is done every 20,000 timesteps for 50 episodes. Both the actor and critic networks contain three hidden layers of size 400, and optimization is done using Adam (Kingma & Ba, 2017). Other algorithm hyperparameters were kept at the implementation’s default values.
@@ C.2 Policy training
As reward model predictions for the starting state can vary across initial states, even for the same task, we shift the rewards of the episode such that $r_{1}=0$, as the subsequent behavior should be scored relative to this state. In addition to this, we apply the standard normalization logic from Gymnasium (Towers et al., 2023), which scales reward values such that their exponential moving average has fixed variance $(1-\gamma)^{2}$.
@@
Training length and hence computational cost varies considerably across tasks. We terminate training after convergence to $\geq$98% success (averaged over the 10 most recent evaluations), or after a maximal training length set per task. The resulting average training length across all 25 tasks was 600k environment steps for VLC MW50 and 750k for the Open X trained model. The corresponding GPU hours vary slightly based on exact architecture used, but we obtained approximately 750k steps in 24 hours on a single NVIDIA V100 or P100 GPU. Our total computational budget was therefore 19.2 GPU hours $\times$ 5 random seeds $\times$ 25 tasks = 2,400 GPU hours for the RL training experiments with VLC MW50 and 24 $\times$ 5 $\times$ 25 = 3,000 hours with VLC Open X. For the MW40 task generalization experiments, the average training length was again 600k for VLC, but 10 random seeds were used, so the corresponding cost for training on 6 tasks was approximately 1150 GPU hours for our method, and slightly more for baselines that took longer to train. Running inference on the VLC architecture to predict the reward for one time step (with batch size 1) takes up an additional 800 MB of GPU RAM, and 11 ms on a H100 GPU or 29 ms on a V100.
Appendix D Meta-World environment definitions
@@
The Meta-World environments are a set of continuous control robotic manipulation tasks. The state space of each environment is $\mathbb{R}^{39}$, containing the positions and orientations of relevant objects as well as the xyz-position of the robot’s end-effector. An action $a$ in $1,1]^{4}$ consists of the desired end-effector xyz-translation, and one dimension for controlling the openness of the gripper. We refer to Yu et al. (2019) for more details on the state information.
As task captions, we use the first sentences of the environment descriptions included in Table 2 in Appendix A - Task Descriptions in Yu et al. (2019). These are fairly succinct and intuitive (examples: Rotate the faucet counter-clockwise, Press a button from the top, Pull a handle up) and we did not need to modify them further.
Appendix E Baseline implementations
@@
For the LOReL baseline experiment, we used the proposed binary classification objective, reversed negatives and 2-frame conditioning (first and last), while keeping all of our other training details and data augmentations identical to our method. This is to ensure LOReL’s smaller and older original architecture as well as lack of visual representation pretraining did not account for any difference in performance. The original implementations were used for all other methods, and pretrained checkpoints (e.g., HowTo100M (Miech et al., 2019) pretraining for RoboCLIP) were reused when applicable. Pretraining datasets, architectures, training objectives and history lengths for each model are listed in Table 6. The difference between temporal aggregation and cosine (or Euclidean) based architectures is illustrated in Figure 5.
@@
Although not considered as a source of data in the original works, we also use our failure videos as additional negatives in the contrastive objective of LIV, RoboCLIP and R3M. Model selection was performed based on validation loss on held-out trajectories. Otherwise, the training procedure and hyperparameter settings were kept unchanged from the original works. In policy training, we apply identical reward normalization (offset and scale, as described in Section 4) for all methods, and reuse the multiplier 50 for the sparse reward.
Table 6: Key differences to prior work. Previous methods use various objectives, but VLC is the first to include a sequence ranking loss, encouraging smooth, dense rewards. Moreover, our model’s conditioning on 12 frames instead of just 1 or 2 like each prior work apart from RoboCLIP allows history-aware progress evaluation, which is important for a variety of tasks (such as distinguishing Turn faucet clockwise and Turn faucet counter-clockwise). Finally, VLC defines rewards based on the output of the temporal Transformer, which is more expressive than compressing the text and video separately into an embedding space and comparing them with either cosine or Euclidean distance, as done by all prior works other than LOReL. As LOReL did not originally use any visual pretraining and used a small architecture, for a fair comparison, we compared to a reimplementation of its training objective while reusing CLIP pretraining and our architecture.
[0.75pt] Objective Num. images Architecture Reward defined by Pretraining data
VLC (ours) CLIP +
sequence ranking
12 CLIP + temporal Transformer model output (temporal Transformer) CLIP
LOReL (original work) binary classification 2 12-layer CNN +
small classifier &
DistilBERT (Sanh, 2019)
model output language only
LOReL (our reimplementation) binary classification 2 CLIP + temporal Transformer model output (temporal Transformer) CLIP
Robo-CLIP CLIP 32 S3D
(Xie et al., 2018)
cosine Howto100M (Miech et al., 2019)
LIV CLIP + VIP
(variant of time-contrastive; Ma et al. (2023b))
1 CLIP
(ResNet-based)
cosine CLIP +
EPIC-KITCHENS (Damen et al., 2018)
Voltron visual reconstruction +
language generation
2 Transformer encoder-decoder &
DistilBERT
Euclidean Something-Something v2 (Goyal et al., 2017)
R3M CLIP + time-contrastive (Sermanet et al., 2018) +
L1 sparsity
1 ResNet (He et al., 2016) & DistilBERT Euclidean Ego4D (Grauman et al., 2022)
[0.75pt]
(a) VLC architecture: temporal aggregation is performed based on both the caption embedding and the sequence of image embeddings.
(b) Distance metric based aggregation: all temporal information is compressed into a fixed-size embedding before being compared with the caption embedding. In addition to the cosine and Euclidean based methods listed in Table 6, the sequence Transformer of Luo et al. (2022) also falls in this category.
(c) Architecture with mean pooling temporal aggregation. Proposed for video retrieval in Luo et al. (2022), but not well suited to reward modeling due to order invariance.
Figure 5: Schematics of the differences between our temporal aggregation architecture (a), architectures that apply a predefined distance metric on top of separate embeddings of the text and the video (b; illustrated with cosine distance, but Euclidean distance has also been used), and mean pooling based architectures (c). Regardless of the choice of text and visual encoders, the main bottleneck in expressiveness in the architectures depicted in b), commonly used in vision-language reward models, comes from compressing the video into a single visual encoding with fixed dimensionality before conditioning on the task caption against which the video should be evaluated, whereas in our architecture, frame-wise information is preserved until the final aggregation step, which is also conditioned on the task caption.
Appendix F Robustness to visual variations
@@
Appendix G Qualitative evaluation
@@
Figure 9: VLC MW40 reward predictions overlaid with example episodes from the Plate Slide Side (above; caption: Slide a plate into a cabinet sideways) and Button Press Top-down Wall tasks (below; caption: Bypass a wall and press a button from the top). The corresponding videos are included on the project website: https://sites.google.com/view/video-language-critic.